Need help with solving an equation
1.06*(1.24/x)= 1.04
Can you please show step by step method to find the value of x. Thank you.

Answer by Alan3354(69443)

1.06*(1.24/x)= 1.04
(1.24/x)= 1.04/1.06
Invert
x/1.24 = 1.06/1.04
x = 1.24*1.06/1.04
=============
Or
1.06*(1.24/x)= 1.04
1.06*1.24= 1.04*x
x = 1.06*1.24/1.04